# Leveraging Self-Supervised Audio-Visual Pretrained Models to Improve Vocoded Speech Intelligibility in Cochlear Implant Simulation

### v2

@@ II-E2 Model Training
We optimized the SSL-AVSE model using partial fine-tuning (PF). Here, the weights pertaining to feature extraction in AV-HuBERT are fixed, whereas the weights in the SSL block are updated based on the pre-trained checkpoint. Training was conducted using an Adam optimizer with a weight decay of $10^{-4}$, a batchsize of 32 and an initial learning rate of $10^{-4}$. In addition, the learning rate is halved when errors are encountered. The proposed model was trained on the TMSV dataset for 50 epochs. We avoid overfitting on the training set by employing early-stopping techniques.
III Experiments
III-A Experimental Setup
@@
Noise can be categorized into static and dynamic based on the acoustic properties in the frequency domain. For example, monotonic background noises would fall under the category of static noise, while highly varied ones would be considered as dynamic. The removing of dynamic noise poses a greater challenge due to the similarities between the input and noise signals such as frequency bands and tonal pitch. To demonstrate the effectiveness of our models, we selected five different dynamic noise types: “babycry” (sound of a baby crying), “babble” (multiple people talking simultaneously in a crowd), “one talker”, “two talkers” and “three talkers”. All these types are associated with human sounds or utterances. The amount of noise used to corrupt the input signal is measured by the signal-to-noise ratio (SNR) and expressed in decibels (dB). The model was trained on the TMSV dataset with video recordings of 18 native Mandarin speakers (13 male and 5 female), each containing 320 utterances. We used 8 speakers (4 male and 4 female) for training and 2 speakers for testing (1 male and 1 female). Each sentence consists of 10 Chinese characters and has a duration of 2 to 4 seconds. The video was recorded with a resolution of 1,920 pixels $\times$ 1,080 pixels at 50 frames per second (fps) while the audio was recorded at a sampling rate of 48 kHz. The first 200 sentences were used to form the training set, while the last 120 sentences were used to form the test sets. To form clean-noisy speech pairs, the training utterances were artificially corrupted by 100 types of noise at five different SNR levels ranging from -12 to 12 dB with an increment of 6 dB. This process generates approximately 600-hour noisy utterances, including the 9-hour randomly selected training data. Instead of using all generated utterances, the training set used in this study comprised 12,000 utterances. Testing was conducted on clean speech mixed with the five dynamic noise types at SNRs ranging from -7 dB to 8 dB at increments of 3 dB.
III-B Experimental Results
III-B1 Effect of Fine-tuning the SSL Pretrained Model
@@
First, we present the speech quality (in terms of PESQ) and speech intelligibility (in terms of STOI) of SSL-AVSE with different pre-trained AV-HuBERT models and verify the effectiveness of fine-tuning these models for AVSE. Table I shows the PESQ and STOI scores of SSL-AVSE using several AV-HuBERT models pre-trained on different datasets, including LRS3 [1] and VoxCeleb2 [10], provided by the authors in [35]. From the table, we can observe that fine-tuning a pre-trained AV-HuBERT model with more diverse data leads to enhanced results. This confirms the effectiveness of our method for leveraging languages with high resources to a use case targeted for another language with low or limited resources. Besides, the results in Table I confirm the efficacy of fine-tuning an AV-HuBERT in our method. In the following experiments, SSL-AVSE denotes the one fine-tuning the AV-HuBERT pre-trained on LRS3, VoxCeleb2 and noise augmentation, which is the best setup.
@@
TABLE I: Objective scores of the speech enhanced by fine-tuning pre-trained AV-HuBERT models. L, V, and N represent LRS3, VoxCeleb2, and noise augmentation, respectively.
@@
III-B2 Objective Results
TABLE II: Objective PESQ scores for non-vocoded speech. We can see that the lower the SNR, the greater the difference between the PESQ scores of SSL-AVSE enhanced audio and those enhanced by either AVSE or AOSE. The converse is true for higher SNRs; the amount of improvement is less marked. However, the higher SNRs, the greater the difference between the PESQ scores of SSL-AVSE enhanced audio and those of the noisy baseline. A stands for audio-only, while AV stands for audio-visual.
@@ III-B2 Objective Results
TABLE III: Objective STOI scores for non-vocoded speech. We can see that the lower the SNR, the greater the difference between the STOI scores of SSL-AVSE enhanced audio and those enhanced by either AVSE or AOSE. For higher SNRs, the results for different enhancement methods converge. A stands for audio-only, while AV stands for audio-visual.
@@
TABLE IV: Objective NCM scores for vocoded speech. Similar to the results of STOI scores, We can see that the lower the SNR, the greater the difference between the NCM scores of SSL-AVSE enhanced audio and those enhanced by either AVSE or AOSE. However, unlike the STOI results, there is no marked convergence between the absolute amount of improvement between the NCM scores of the noisy baseline and that of utterances enhanced by SSL-AVSE.
@@
As shown in Tables II to IV, compared with AOSE and AVSE, enhancement by SSL-AVSE resulted in higher objective measures of speech intelligibility for most types of noise. The difference was particularly notable for low SNRs. For a SNR of -7 dB, the PESQ, STOI and NCM values of SSL-AVSE were 3.7% (from 1.226 to 1.271), 6.0% (from 0.547 to 0.582) and 9.3% (from 0.387 to 0.423) higher than those of AVSE, respectively, while they were 3.6% (from 1.227 to 1.271), 11.3% (from 0.523 to 0.582) and 19.5% (from 0.354 to 0.423) higher than those of AOSE, respectively. The results were even more striking when compared with the noisy baseline; PESQ, STOI and NCM values increased by 4.7% (from 1.214 to 1.271), 8.6% (from 0.536 to 0.582) and 87.2% (from 0.226 to 0.423), respectively. However, for higher SNRs, the differences between the PESQ, STOI and NCM values of SSL-AVSE and AVSE enhanced speech are less pronounced; this is probably because automatic labelling is less useful when fewer distinguishing features are present. The presence of visual cues were also found to be less effective at audio denoising for higher SNRs; results also show reduced improvement in all three speech quality scores in audio enhanced by SSL-AVSE over those enhanced by AOSE.
TABLE V: Paired t-test for speech quality under selected noise conditions. The degree of freedom for all values is 19. A p-value of less than 0.05 implies statistical significance.
@@
AOSE 1.571 $<$0.05 1.746 $<$0.05 1.191 $<$0.05 3.145 $<$0.05
AVSE 1.337 $<$0.01 0.883 $<$0.01 1.047 $<$0.01 2.732 $<$0.05
SSL-AVSE 2.743 $<$0.001 1.402 $<$0.001 1.031 $<$0.001 1.433 $<$0.01
TABLE VI: Paired t-test for speech intelligibility under selected noise conditions. The degree of freedom for all values is 19. A p-value of less than 0.05 implies statistical significance.
@@
AOSE 2.518 $<$0.05 3.452 $<$0.05 2.346 $<$0.05 2.128 $<$0.05
AVSE 1.664 $<$0.01 3.683 $<$0.05 2.038 $<$0.05 3.038 $<$0.01
SSL-AVSE 2.842 $<$0.001 4.535 $<$0.01 1.406 $<$0.01 4.406 $<$0.01
III-B3 Spectrogram Analysis
A spectrogram plot is frequently employed to visually represents the time-frequency characteristics of a speech signal. In Fig. 4, we present spectra of a noisy speech signal at a 2 dB SNR, enhanced using four methods: AOSE, ConvTasNet, AVSE, and SSL-AVSE. Additionally, the spectra of the corresponding clean speech is included for comparison. It can be observed that SSL-AVSE outperforms AOSE, ConvTasNet and AVSE in effectively removing noise components and restoring the original clean speech. In Fig. 5, we showcase spectra of vocoded speech. From the figure, it is evident that the spectra of vocoded speech processed by SSL-AVSE preserves much clearer speech structures compared with AOSE and AVSE.
Figure 4: Spectrograms of clean, noisy, AOSE enhanced, AVSE enhanced, ConvTasNeT enhanced and SSL-AVSE enhanced audio signals for “babble 2 dB” noise. Note that SSL-AVSE enhanced audio emphasizes audio tones within the range of human speech more than those of other enhancement methods.
Figure 5: Spectrograms of vocoded speech for noisy, clean, AOSE enhanced, AVSE enhanced and SSL-AVSE enhanced audio signals for “babble 2 dB” noise. Note that SSL-AVSE-enhanced speech has clearer structures compared with those of other enhancement methods.
III-B4 Subjective Results
We conducted listening tests using four different types of noises. Each noise types consisted 120 utterances, with 20 from each of the following six categories: AOSE, AVSE, SSL-AVSE, clean, logarithm minimum mean squared error (logMMSE), and noisy speech. The logMMSE method performs SE based on assumptions of distribution and properties of speech and noise signals[15]. Similar to[50], we also include logMMSE as part of our benchmark. In this study, we also selected to present those utterances corrupted by “babycry” and “babble” noises with SNRs of 2 dB and 5 dB, respectively. Since that our objective is to maximize SE in CI devices, we conducted our experiment using vocoded speech. A total of 80 participants with a mean age of 35.7 took part in the study, with 20 individuals assigned to each noise or sound condition to reduce cross-referencing bias. The listening tests were conducted in a quiet room, while the utterances were uploaded onto a specially designed system that presented them randomly to the users.
@@ III-B4 Subjective Results
Figure 6: Subjective sound quality scores for vocoded speech enhanced with different models. The x-axis represents the noise type while the y-axis represents the sound quality score. For all noise types, SSL-AVSE is shown to perform notably better than other methods, with the greatest improvements occurring for noises with an SNR of 2 dB. logMMSE, a statistical SE method, is shown to perform worse than the noisy baseline.
@@
Figure 7: Subjective character intelligibility scores of speech enhanced using different models. The x-axis represents the noise type while the y-axis represents the word intelligibility score. As in the case of speech quality, SSL-AVSE is also shown to perform notably better than other methods, with the greatest improvements occurring for noises with an SNR of 2 dB.
@@
Figure 8: Relationship between the subjective and objective scores of speech enhanced using different models. We choose utterances corrupted with “babble 2 dB” noise. The x-axis represents the subjective score while the y-axis represents the objective score. The centers of the ovals represent the mean objective and subjective scores, while the lengths of the major and minor axes represent the standard deviations of the objective and subjective scores, respectively. There is clear segregation between utterances enhanced using the three different methods, with SSL-AVSE having the best effect.
@@
We further conducted paired t-tests on the quality and intelligibility results of listening test under SSL-AVSE, AVSE and AOSE conditions to verify their statistical significance. We specifically compare the outcomes obtained from “babble” and “babycry” noises for SNRs of 2dB and 5dB, since they are generally considered the most difficult to denoise. Similar trends are observed for other types of noise as well. Results shown in Tables V to VI indicate that audio enhanced by all three methods differ statistically from noisy audio, with $p$-values all less than 0.001 for “babble” noises and $p$-values all less than 0.05 for “babycry” noises. Since the “babble” noise condition results in smaller $p$-values than those of “babycry”, this demonstrates that our model actually performs better under more challenging noise conditions, especially those with multiple talkers. Further comparisons between audio enhanced by SSL-AVSE and those enhanced by either AVSE or AOSE also show statistical differences, further confirming the utility of our proposed models.
III-B5 Comparison between Objective and Subjective Results
@@
Both objective and subjective evaluations demonstrate that the performance measures improve with SNR levels. Moreover, it is observed that audio signals with lower SNRs exhibit a more substantial enhancement in speech quality, highlighting the robustness of SSL-AVSE in challenging noisy environments. The subjective results further validate the effectiveness of our model in handling vocoded speech, showing improvements comparable to those determined by objective methods when compared to the noisy baseline. To more explicitly explore the relationship between objective and subjective results, three panels are presented in Fig. 8, focusing on utterances corrupted with “babble 2 dB” noise. For all three panels, the x-axis represents the subjective score, while the y-axis represents the objective score. Analyzing the left panel of Fig. 8, it is evident that SSL-AVSE outperforms AVSE and AOSE notably in terms of both objective and subjective speech quality scores, as indicated by the score distributions located towards the top-right side. Furthermore, examining the center and right panels of Fig. 8, it is evident that SSL-AVSE achieves the best performance for both objective and subjective speech intelligibility scores, as indicated by the score distributions also located towards the top-right side.
IV Discussion
In this paper, we propose a deep neural network-based AVSE method that leverages both audio and visual signals, evaluating its effectiveness in enhancing speech perception in CI simulations. As shown in Table IV, when combined with self-supervised learning, NCM values of enhanced audio increased by between 26.5% (from 0.671 to 0.846 for 8dB) and 87.2% (from 0.226 to 0.423 for -7dB) compared with the noisy baseline. This demonstrates the effectiveness of our model on vocoded audio signals. Subjective results also showed improvements of between 12.3% and 45.2% for sound quality (from 3.59 to 4.30 for “babycry 5dB” and from 2.10 to 3.05 for “babble 2dB”) and improvements of between 5.6% and 51.8% for word intelligibility (from 0.654 to 0.691 for “babycry 2dB” and from 0.372 to 0.565 for “babble 2dB”) when compared with the noisy baseline, confirming our results.
Due to the relatively large size of video files, real-time AVSE using both video and audio signals remains a challenging task. First, an appropriate image extraction rate needs to be determined based on trade-offs between testing time and accuracy. Second, most SSL models are relatively large; for example, our base AV-HuBERT model contains more than 300 million parameters. Reducing the number of parameters through prudent data pre-processing or by pruning the model to reduce its size after training yield faster results at the expense of testing accuracy. For settings that require high fidelity output audio including lectures and meetings, this may not be feasible. Further work is required to determine the optimal parameters for each setting. While our current model may not directly suit CI devices with limited capacity, it can be implemented on external interface hardware like smartphones, providing ample computational resources for speech signal processing. For example, CIs have started to connect with phones as a “router” such as in the Cochlear Wireless PhoneClip, which may improve speech recognition by converting Bluetooth sound streams to the Nucleus 6 (CP910) sound processor [40]. Similarly, the AudioLink device can help with auditory training on speech recognition with compatible MED-EL sound processors [66]. More straightforward designs, such as the Nucleus 7 (CP1000) and more recent sound processors may directly connect to compatible Apple (Made for iPhone, MFi) and Android (Android Streaming for Hearing Aids, ASHA) devices [83]. Our system is also well-suited for applications like teleconferencing and playback applications, where real-time processing is not essential. With rapid advancements in hardware, complex systems based on DNNs can now run on edge devices, suggesting that hardware capable of supporting our system will soon be widely available.
V Conclusion
We propose a novel AVSE model that leverages a pretrained audio-visual foundation model to enhance speech intelligibility and quality in CI scenarios. Although the model was pretrained using an English corpus, it performed well on SE tasks involving Mandarin datasets, demonstrating the ability to generalize to different target languages. Compared with other state-of-the-art AVSE methods, our proposed model resulted in a notable increase in speech quality and word intelligibility. In the future, we aim to conduct real-time signal processing by compressing our model using pruning and data quantization techniques. This enables more rapid processing of audio-visual signals without compromising significantly on quality. The results can then be output as audio signals or as text files for the patient, enabling them to regain near-normal hearing function.

### v3

@@ 2.5.2 Model Training
We optimized the SSL-AVSE model using partial fine-tuning (PF). In this setup, the convolutional layer weights in AV-HuBERT, used for feature extraction, are kept fixed, while the transformer encoder weights are fine-tuned from the pretrained checkpoint. Training was conducted using the Adam optimizer with a weight decay of $10^{-4}$, a batchsize of 32, and an initial learning rate of $10^{-4}$. In addition, the learning rate is halved when errors are encountered. The proposed model was trained on the Taiwan Mandarin speech with video (TMSV) dataset  [^1] https://bio-asplab.citi.sinica.edu.tw/Opensource.html#TMSV for 50 epochs. We avoid overfitting on the training set by employing early-stopping techniques.
3 Experiments
3.1 Experimental Setup
@@
Noise can be roughly categorized into stationary and non-stationary based on the acoustic properties in the frequency domain. For example, monotonic background noises would fall under the category of stationary noise, while highly varied ones would be considered non-stationary. The removal of non-stationary noise is particularly challenging due to the substantial overlap and interference between speech and noise signals. To assess the effectiveness of our models, five non-stationary noise conditions were considered in the test set: “babycry” (sound of a baby crying), “babble” (multiple people talking simultaneously in a crowd), “one talker”, “two talkers”, and “three talkers”. All these types are associated with human sounds or utterances. The amount of noise used to corrupt the input signal is measured by the signal-to-noise ratio (SNR) and expressed in decibels (dB).
The SE models were trained on the TMSV dataset, which includes video recordings of 18 native Mandarin speakers (13 male and 5 female), each contributing 320 utterances. Eight speakers (four male and four female) were used for training, while an additional unseen male speaker was reserved for testing. Each utterance consists of 10 Chinese characters and lasts between 2 to 4 seconds. The video was recorded with a resolution of 1,920 pixels $\times$ 1,080 pixels at 50 fps while the audio was recorded at a sampling rate of 48 KHz. The first 200 sentences were used to form the training set, while the last 120 sentences were used to form the test set. To create noisy–clean speech pairs, the training utterances were artificially corrupted with 100 types of noise at five SNR levels ranging from –12 to 12 dB in 6 dB increments, yielding approximately 600 hours of noisy speech. Instead of using all generated utterances, the training set was constructed from 12,000 randomly selected noisy–clean pairs drawn from the full dataset. Testing was conducted on clean speech mixed with the five non-stationary noise types as mentioned earlier at SNRs ranging from -7 dB to 8 dB at increments of 3 dB.
3.2 Experimental Results
3.2.1 Effect of Fine-tuning the SSL Pretrained Model
@@
First, we present the speech quality (in terms of PESQ) and speech intelligibility (in terms of STOI) of SSL-AVSE with different pretrained AV-HuBERT models and verify the effectiveness of fine-tuning these models for AVSE. Table 1 shows the PESQ and STOI scores of SSL-AVSE using several AV-HuBERT models pretrained on different datasets, including LRS3 [101] and VoxCeleb2 [102], and noise augmentation, provided by the authors in [75]. From the table, we can observe that fine-tuning a pretrained AV-HuBERT model with more diverse data leads to enhanced results. Since both LRS3 and VoxCeleb2 are English datasets, while the testing data consists of Mandarin speech, the results in Table 1 demonstrate the potential of the proposed SSL-AVSE method to leverage high-resource languages for applications in low-resource languages. Furthermore, the results in Table 1 confirm the effectiveness of fine-tuning AV-HuBERT within our method. In the following experiments, SSL-AVSE denotes the one fine-tuning the AV-HuBERT pretrained on LRS3, VoxCeleb2, and noise augmentation, which is the best setup in Table 1.
@@
Table 1: Objective scores of the speech enhanced by fine-tuning pretrained AV-HuBERT models. L, V, and N represent LRS3, VoxCeleb2, and noise augmentation, respectively, and w/o denotes without fine-tuning the AV-HuBERT model.
@@
3.2.2 Objective Results
Table 2: Objective PESQ scores for non-vocoded speech. We can see that the lower the SNR, the greater the difference between the PESQ scores of SSL-AVSE-enhanced speech and those enhanced by either AVSE or AOSE. At higher SNRs, the differences between enhancement methods become less pronounced. ”A” stands for audio-only, while ”AV” stands for audio-visual.
@@ 3.2.2 Objective Results
Table 3: Objective STOI scores for non-vocoded speech. We can see that the lower the SNR, the greater the difference between the STOI scores of SSL-AVSE-enhanced speech and those enhanced by either AVSE or AOSE. For higher SNRs, the results for different enhancement methods converge. ”A” stands for audio-only, while ”AV” stands for audio-visual.
@@
Table 4: Objective LPS scores for non-vocoded speech. We can see that the lower the SNR, the greater the difference between the LPS scores of SSL-AVSE-enhanced speech and those enhanced by either AVSE or AOSE. At higher SNRs, the differences between enhancement methods become less pronounced. ”A” stands for audio-only, while ”AV” stands for audio-visual.
modality -7dB -4dB -1dB 2dB 5dB 8dB
Noisy N/A 0.146 0.187 0.244 0.322 0.417 0.528
AOSE A 0.196 0.235 0.281 0.349 0.444 0.555
AVSE-VAE [105] AV 0.167 0.197 0.246 0.322 0.405 0.495
AVSE AV 0.204 0.254 0.311 0.38 0.483 0.587
SSL-AVSE AV 0.275 0.32 0.387 0.463 0.559 0.639
Table 5: Objective NCM scores for vocoded speech. Similar to the results of STOI scores, We can see that the lower the SNR, the greater the difference between the NCM scores of SSL-AVSE-enhanced speech and those enhanced by either AVSE or AOSE. However, unlike the STOI results, there is no marked convergence between the absolute amount of improvement between the NCM scores of the noisy baseline and that of utterances enhanced by SSL-AVSE.
@@
As shown in Tables 25, the proposed SSL-AVSE consistently outperformed the baseline AOSE and AVSE systems, achieving higher objective measures of speech quality, intelligibility, and LPS across most noise conditions. The difference was particularly notable for low SNRs. For an SNR of -7 dB, the PESQ, STOI, LPS and NCM values of SSL-AVSE were 3.7% (from 1.226 to 1.271), 6.0% (from 0.549 to 0.582), 34.8% (from 0.204 to 0.275), and 9.3% (from 0.387 to 0.423) higher than those of AVSE, respectively, while they were 3.6% (from 1.227 to 1.271), 11.3% (from 0.523 to 0.582), 40.3% (from 0.196 to 0.275), and 19.5% (from 0.354 to 0.423) higher than those of AOSE, respectively. The results were even more striking when compared with the noisy baseline; PESQ, STOI, LPS, and NCM values increased by 4.7% (from 1.214 to 1.271), 8.6% (from 0.536 to 0.582), 88.4% (from 0.146 to 0.275), and 87.2% (from 0.226 to 0.423), respectively. At higher SNRs, the differences in PESQ, STOI, and NCM scores between SSL-AVSE and AVSE are less pronounced, indicating that the pretrained AV-HuBERT model provides greater benefits for SE under challenging conditions. The results also show smaller improvements across all three metrics for speech enhanced by SSL-AVSE compared to AOSE at higher SNR levels, indicating that visual cues provide relatively less benefit for AVSE performance under these conditions.
Table 6: Paired t-tests of speech quality results were used to compare AOSE, AVSE, and SSL-AVSE against the noisy baseline for each specific noise type and SNR condition. The degree of freedom for all values is 19. A p-value of less than 0.05 implies statistical significance.
@@
AOSE 1.571 <<0.05 1.746 <<0.05 1.191 <<0.05 3.145 <<0.05
AVSE 1.337 <<0.01 0.883 <<0.01 1.047 <<0.01 2.732 <<0.05
SSL-AVSE 2.743 <<0.001 1.402 <<0.001 1.031 <<0.001 1.433 <<0.01
Table 7: Paired t-tests of speech intelligibility results were used to compare AOSE, AVSE, and SSL-AVSE against the noisy baseline for each specific noise type and SNR condition. The degree of freedom for all values is 19. A p-value of less than 0.05 implies statistical significance.
@@
AOSE 2.518 <<0.05 3.452 <<0.05 2.346 <<0.05 2.128 <<0.05
AVSE 1.664 <<0.01 3.683 <<0.05 2.038 <<0.05 3.038 <<0.01
SSL-AVSE 2.842 <<0.001 4.535 <<0.01 1.406 <<0.01 4.406 <<0.01
3.2.3 Spectrogram Analysis
A spectrogram plot is frequently employed to visually represent the time–frequency characteristics of a speech signal. In Fig. 4, we present spectrograms of a noisy speech signal at a 2 dB SNR, enhanced using four methods: AVSE-VAE [105], AOSE, AVSE, and SSL-AVSE. Additionally, the spectra of the corresponding clean speech are included for comparison. Two regions of interest are highlighted in the spectrograms: a noise-only segment (yellow box) and a mixed speech–noise region (green dashed box). In the noise-only regions (yellow box), SSL-AVSE shows significant improvements over the baseline methods by effectively reducing noise. In the mixed speech–noise regions (green dashed box), SSL-AVSE introduces fewer distortions in the reconstructed speech compared with the other approaches. In Fig. 5, we showcase spectra of vocoded speech. From the figure, it is evident that the spectra of vocoded speech processed by SSL-AVSE preserve much clearer speech structures compared with AOSE and AVSE.
Figure 4: Spectrograms of noisy, clean, AVSE-VAE enhanced, AOSE enhanced, AVSE enhanced, and SSL-AVSE-enhanced speech signals for “babycry 2 dB” noise. Note that SSL-AVSE enhanced speech preserves speech structures within the range of human speech more than those of other enhancement methods.
Figure 5: Spectrograms of vocoded speech for noisy, clean, AOSE enhanced, AVSE enhanced, and SSL-AVSE-enhanced speech signals for “babycry 2 dB” noise. Note that SSL-AVSE enhanced speech has clearer structures compared with those of other enhancement methods.
3.2.4 Subjective Results
We conducted listening tests under four conditions, comprising two noise types (“babycry” and “babble”) and two SNRs (2 dB and 5 dB). Each condition contained 120 utterances, with 20 drawn from each of six categories: AOSE, AVSE, SSL-AVSE, clean, logarithm minimum mean squared error (logMMSE), and noisy speech. The logMMSE is a traditional SE method that enhances speech by minimizing the mean-square error in the logarithmic spectral domain [106]. Similar to [80], we also include logMMSE as part of our benchmark. Since our objective is to maximize SE in CI devices, we conducted our experiment using vocoded speech. A total of 80 participants (mean age: 35.7 years) took part in the study, with 20 individuals assigned to each noise type at a specific SNR condition to minimize cross-referencing bias. The listening tests were conducted in a quiet room, while the utterances were uploaded onto a listening test system that presented them randomly to the users.
@@ 3.2.4 Subjective Results
Figure 6: Subjective speech quality scores for vocoded speech enhanced with different models. The x-axis represents the noise type, while the y-axis represents the speech quality score. For all noise types, SSL-AVSE is shown to perform notably better than other methods, with the greatest improvements occurring for noises with an SNR of 2 dB.
@@
Figure 7: Subjective character intelligibility scores of speech enhanced using different models. The x-axis represents the noise type, while the y-axis represents the word intelligibility score. As in the case of speech quality, SSL-AVSE is also shown to perform notably better than other methods, with the greatest improvements occurring for noises with an SNR of 2 dB.
@@
Figure 8: Relationship between the subjective and objective scores of speech enhanced using different models. We choose utterances corrupted with “babble 2 dB” noise. The y-axis represents the subjective score, while the x-axis represents the objective score. The centers of the ovals represent the mean objective and subjective scores. There is clear segregation between utterances enhanced using the three different methods, with SSL-AVSE having the best effect.
@@
We further conducted paired t-tests on the quality and intelligibility results of listening tests under SSL-AVSE, AVSE, and AOSE conditions to verify their statistical significance. Results shown in Tables 6 and 7 indicate that speech enhanced by all three methods differ statistically from noisy speech, with $p$-values all less than 0.001 for the “babble” noise and $p$-values all less than 0.05 for the “babycry” noise. Since the “babble” noise condition results in smaller $p$-values than those of “babycry”, this demonstrates that our model actually performs better under more challenging noise conditions, especially those with multiple talkers. Further comparisons between speech enhanced by SSL-AVSE and that enhanced by AVSE or AOSE revealed significant statistical differences, reinforcing the promising capability of the proposed method.
3.2.5 Comparison between Objective and Subjective Results
@@
Both objective and subjective evaluations demonstrated that SSL-AVSE yielded improved performance across different SNR levels. Moreover, speech signals with lower SNRs exhibited greater improvements in quality and intelligibility, underscoring the strong capability of SSL-AVSE in challenging noisy environments. The subjective results further validate the effectiveness of our model for vocoded speech, showing notable improvements compared with the noisy baseline. To more explicitly explore the relationship between objective and subjective results, three panels are presented in Fig. 8, focusing on utterances corrupted with “babble 2 dB” noise. For all three panels, the x-axis represents the subjective score, while the y-axis represents the objective score. Analyzing the left panel of Fig. 8, it is evident that SSL-AVSE outperforms AVSE and AOSE notably in terms of both objective and subjective speech quality scores, as indicated by the score distributions located towards the top-right side. Furthermore, examining the center and right panels of Fig. 8, it is evident that SSL-AVSE achieves the best performance for both objective and subjective speech intelligibility scores, as indicated by the score distributions also located towards the top-right side.
4 Discussion
[…]
Our experimental results also indicate that the proposed SSL-AVSE model, although leveraging an audio-visual foundation model pretrained on English corpora, can still achieve strong performance on Mandarin AVSE tasks. We believe that this cross-lingual transferability is partly due to two factors. First, regression tasks such as speech enhancement and separation tend to be less language-dependent because they primarily focus on improving acoustic properties rather than modeling linguistic content. Prior work has shown that models trained exclusively on English can generalize well to other languages (e.g., Spanish and German) for SE without additional adaptation [107]. Second, audio-visual alignment in models like AV-HuBERT is largely based on shared phoneme-level features, which are relatively consistent across languages. This allows the learned correlations between phonemes and visual articulations (e.g., lip shapes) to transfer more effectively. Recent studies [68, 108] have shown that English-pretrained models (e.g., AV-HuBERT, U-Net) can improve Mandarin AVSE performance even without retraining the visual front-end, whereas word-level tasks such as lip reading often require language-specific mechanisms [109]. These observations suggest why the SSL-AVSE system can generalize effectively to Mandarin tasks despite being pretrained on English corpora.
Table 8: Model complexity and inference latency of the proposed SSL-AVSE system.
Metric Value
Parameters 106 M (103 M for AV-HuBERT)
FLOPs 27.5 GFLOPs
Model Size 423 MB (FP32) / 107 MB (INT8)
Latency – 5 ms per 10 ms audio frame (measured on NVIDIA GeForce RTX 2080 Ti)
– 66 ms per 10 ms audio frame (measured on 4-core Intel Xeon Gold 6152 CPU, ONNX Runtime)
Table 9: Performance comparison of different model setups before and after 8-bit quantization. WF, PF, and EF denote without fine-tuning, partial fine-tuning, and entire fine-tuning, respectively, on AV-HuBERT. Quantization introduces only minor degradation in PESQ, while STOI scores slightly improve across all SSL-based models.
PESQ(Orig) PESQ (8-bit) $\Delta$ PESQ STOI (Orig) STOI (8-bit) $\Delta$ STOI
AVSE 1.245 1.237 -0.009 0.605 0.604 -0.001
SSL-AVSE (WF) 1.299 1.253 -0.046 0.630 0.641 +0.011
SSL-AVSE (EF) 1.374 1.339 -0.036 0.661 0.664 +0.003
SSL-AVSE (PF) 1.396 1.361 -0.035 0.682 0.686 +0.004
It is important to note that the performance of our system was evaluated using a vocoder-based CI simulation framework. In CI research, vocoder-based CI simulation has become a widely used method, as supported by prior studies [88, 110, 91]. Although such simulations cannot fully capture the auditory experience of actual CI users, they have been shown to reproduce similar behavioral performance trends. This approach offers several key advantages. First, it reduces variability caused by individual factors in CI users, such as etiology of hearing loss, duration of deafness, and electrode placement. Second, it reduces the fatigue and discomfort that CI users may experience during extended testing sessions, thereby preventing inaccurate or biased results. Third, using vocoded speech with normal-hearing participants allows researchers to isolate and evaluate the effects of signal processing algorithms without confounds from CI hardware differences or user-specific neural adaptation. Given recent advances in CI technology, many studies now use 16-channel tone vocoders [111, 112]. Accordingly, we adopt the same configuration in this study.
To evaluate the feasibility of deploying the proposed SSL-AVSE system in real-world applications, we conducted a detailed analysis of its model complexity and computational efficiency. As summarized in Table 8, the SSL-AVSE system comprises 106 million (M) parameters (103 M for the AV-HuBERT encoder) and requires 27.5 GFLOPs per inference. Profiling results further indicate that the model achieves 5 ms processing time per 10 ms audio frame on the GPU and 66 ms per frame on a 4-core CPU. In addition, we investigated the impact of post-training 8-bit quantization, which reduced the overall model size from 423 MB to 107 MB. As shown in Table 9, quantization introduced only negligible degradation in PESQ, while STOI scores improved slightly across all SSL-based models. We hypothesize that this effect may stem from quantization discarding non-essential variations, thereby sharpening intelligibility-related representations.
These findings collectively demonstrate that the quantized SSL-AVSE system is already well-suited for deployment on more powerful external computing platforms, such as personal computers, smartphones, smart glasses, or tablets, which can readily accommodate the required computational resources. In the context of CI, such devices could act as intermediate processing units to deliver enhanced speech signals to CI sound processors. Existing commercial solutions already support this type of integration. For instance, the Cochlear Wireless PhoneClip streams Bluetooth audio to the Nucleus 6 (CP910) sound processor [113], whereas AudioLink serves as a universal wireless streamer, transmitting audio from external devices (e.g., phones, tablets, TVs, and remote microphones) directly to MED-EL sound processors [114]. More recent CI processors, such as the Nucleus 7 (CP1000), are capable of directly interfacing with compatible Apple devices (Made for iPhone, MFi) and Android devices (Android Streaming for Hearing Aids, ASHA) [115]. These established integration pathways suggest that the proposed system could feasibly be incorporated into current CI user ecosystems in the near term.
Looking ahead, further model compression strategies, including pruning and knowledge distillation, are expected to significantly reduce the computational cost and memory footprint of the SSL-AVSE system. These advances would open the possibility of directly deploying the system on CI processors with stringent resource constraints, thereby eliminating the reliance on external edge devices. Ultimately, such developments would extend the benefits of advanced SE technologies to CI users in a wider range of real-world acoustic environments.
5 Conclusion
We propose a novel AVSE model, SSL-AVSE, which leverages a pretrained audio-visual foundation model to enhance speech quality and intelligibility in CI scenarios. Although the model was pretrained using an English corpus, it performed well on SE tasks involving Mandarin datasets, demonstrating the ability to generalize to different target languages. Compared with other state-of-the-art AVSE methods, our proposed model resulted in a notable increase in speech quality and word intelligibility. Although many previous studies have demonstrated the effectiveness of pretrained models on various downstream tasks, this work is the first to apply a pretrained audio-visual model to enhance SE performance and demonstrate its potential benefits for CI users. We believe that this study represents a promising direction for advancing AVSE technologies across a range of assistive listening devices, including hearing aids, CIs, and personal sound amplification products (PSAPs). Moving forward, we aim to reduce the model size of the proposed SSL-AVSE system through techniques such as parameter pruning and knowledge distillation. As modern devices increasingly incorporate sensors to capture multimodal data, we also plan to explore the integration of additional modalities, such as tactile and textual information, to further improve SE performance. This represents another important avenue for future research.